OpenStudy (anonymous):

What is the probability of getting atleast one 6 in 3 attempts of a dice?

OpenStudy (anonymous):

1/6 X 1/6 X 1/6 = 1/216 ?

OpenStudy (anonymous):

i feel like it should be better than 1/6... but each roll is independent =/

OpenStudy (anonymous):

I got my question's answer myself 1-(5/6)^3 = 91/256